OpenStudy (anonymous):

find the absolutemaximum and minimum values of each function on the given interval.

OpenStudy (anonymous):

\[f(\theta)=\sin (\theta), -\Pi/2\le \theta \le5\Pi/6\]

OpenStudy (anonymous):

Well, first off all we need to determine where the orginal function i undefined. In this case it is contious on that given interval. Now we have to obtains the derivaitve of the orginal function and determine where it is zero

OpenStudy (anonymous):

Then graph the function. Identify the points on the graph where the absolute extreme occurs, and include their coordinates.

OpenStudy (anonymous):

Now the derivaitve of the sinx function i cos(x). Where is this zero?

OpenStudy (anonymous):

Specifically within the given interval

OpenStudy (anonymous):

the zero for cos?

OpenStudy (anonymous):

yes, where is cos(x) zero? In your given interval, it should b zero at , -pi/2 and pi/2

OpenStudy (anonymous):

thats what i thought but in the solutions manual it says that -pi/2 is not interior to the domain.

OpenStudy (anonymous):

i dont know what they mean by that

OpenStudy (anonymous):

but you said, our domain is -pi/2 to 5pi/6 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so, now, what you will have to do is evalauate each of the points we have to work with, namely, the two endpoints, -pi/2 and 5pi/6, as well as -pi/2 and pi/2

OpenStudy (anonymous):

and you have to evaluate these critical numbers using the orginal function

OpenStudy (anonymous):

Now the highest max, you get is your absolute maximum, and the lowest min you get is your absolute minimum. The rest are relative extrema

OpenStudy (anonymous):

i know how to do everything but im just trying to figure out why i cant use -pi/2; why it says that it doesnt work because it isnt interioir to the domain

OpenStudy (anonymous):

are you sure your boundaries are right, that i to say your interval

OpenStudy (anonymous):

perhaps it is -pi/6

OpenStudy (anonymous):

yeah im positive, thats why im so confused

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

it says it isnt interior to the domain which mean it is part of the domain, but not interior

OpenStudy (anonymous):

Right, -pi/2 is not inside the domain, it is simply the endpoint

OpenStudy (anonymous):

the problem before this one i did the same thing and i used an endpoint as well though

OpenStudy (anonymous):

g(x)=\[\sqrt{4-x ^{2}} -2\le x \le 1\]

OpenStudy (anonymous):

the answers should be pi/2 and -pi/2

OpenStudy (anonymous):

it lets me use -2 but not 2 because it isnt in the domain

OpenStudy (anonymous):

yes, but in out case we can use -pi/2 because it is an endpoint and we have to evaluate the orginal function at that value

OpenStudy (anonymous):

g(x)= so what about in the other one i get a -2 as a critical point but it doesnt tell me i cant use it and its an endpoint as well